Suppose a random sample of 25 measurements is taken from a population with mean 17 and variance 100. The variance of the sample mean is
Answer #1

Solution:

Variance of sample mean = σ^2/n

WE are given

σ^2 =100

n = 25

Variance of sample mean = 100/25 = 4

The variance of the sample mean is 4.